OpenStudy (anonymous):

If P(x)=5x-4/x+3, find P(-4)

OpenStudy (anonymous):

(5*-4)-(4/-4)+3

OpenStudy (anonymous):

replace x with -4 and do the math.

OpenStudy (anonymous):

so my answer would be -7

OpenStudy (anonymous):

please show your work.

OpenStudy (anonymous):

5.4=20-4=5 -4+3=7

OpenStudy (anonymous):

no its suppose to be -4 not 4 that you replace

OpenStudy (anonymous):

try it again using -4

OpenStudy (anonymous):

-24/-1=24

OpenStudy (anonymous):

yes you are correct.

OpenStudy (anonymous):

thanks